What is snell law.
1
Expert's answer
2020-12-25T15:21:25-0500

Snell's law determines the direction of propagation of a ray of light that falls on the interface between two different isotropic media


"\frac{\sin\alpha_2}{\sin\alpha_1}=\frac{v_2}{v_1}=\frac{n_1}{n_2}".